# Learning in Stackelberg Mean Field Games: A Non-Asymptotic Analysis

### v1

@@ 1 Introduction
Mean field games (MFGs) provide a framework for studying the strategic interaction among an infinite number of rational agents, with a wide range of applications in resource allocation (Li et al., 2020), telecommunication (Narasimha et al., 2019), and power system optimization (Alasseur et al., 2020). An important extension of MFGs is Stackelberg mean field games (SMFGs), incorporating a hierarchical structure where a single leader agent influences a population of follower agents and enjoys a first-mover advantage. A prominent example of SMFG is optimal liquidation (Almgren and Chriss, 1997), in which a large institutional investor (leader) strategically sells assets, e.g., central bank treasuries, in a market while accounting for the reactions of smaller traders (followers) who adjust their behavior, e.g., bond purchasing, in response (Chen et al., 2024). Other examples arise in public policy domains such as taxation, mortgage regulation, and epidemic control (Zheng et al., 2020; Mi et al., 2024; Aurell et al., 2022), where the leader, typically a government, designs interventions to promote social welfare while anticipating and influencing the collective response of the population.
Despite the increasing number of empirical studies applying the SMFG framework to real-world problems, there remains a lack of practical and theoretically grounded algorithms for learning in such hierarchical environments. The central challenge lies in the inherently coupled dynamics among the leader, a representative follower, and the mean field representing the collective behavior of the follower population. This interdependence gives rise to a complex bi-level structure: the leader’s objective depends on the equilibrium response of the follower population, which in turn is shaped by the leader’s policy. Existing approaches (Dayanıklı and Laurière, 2024; Cui et al., 2024) to SMFGs (and the closely related major-minor MFGs) need to make strong assumptions to decouple the leader and followers, lack non-asymptotic guarantees for finding a solution, and empirically exhibit slow convergence or tend towards sub-optimal limit points, possibly due to cyclic behaviors. In contrast, our work relaxes the assumptions and establishes a fast finite-sample rate of convergence for a simple actor-critic algorithm to the stationary point of the SMFG objective.
Figure 1: Theoretical convergence rates. The proposed algorithm is the first to have non-asymptotic convergence guarantees (black). Existing algorithms in similar settings are only known to converge asymptotically (red). Convergence rate of a single-loop algorithm for bi-level optimization (Hong et al., 2023) under a non-convex upper-level objective and lower-level strong convexity (blue).
@@ 1.1 Main Contributions
Notably, our convergence rate surpasses that of a comparable single-loop algorithm for bi-level optimization under lower-level strong convexity (Hong et al., 2023) ($\widetilde{{\cal O}}(k^{-1/2})$ vs $\widetilde{{\cal O}}(k^{-2/5})$). The policy optimization problem in a SMFG can be viewed as bi-level optimization, where the upper-level objective is non-convex and the lower-level problem is to solve a standard MFG that does not exhibit any convexity. The reason that we achieve a better rate in a more challenging setting is two-fold. First, as a key technical innovation, we develop an analytical argument to handle bi-level optimization with lower-level Polyak-Lojasiewicz (PL) condition, which the MFG objective satisfies with respect to the follower’s softmax policy parameter under regularization (Mei et al., 2020). This technique allows us to obtain the same convergence rate in SMFGs as in bi-level optimization with lower-level strong convexity, and may be of independent interest and applicable to general bi-level optimization methods. Second, we make the rate improvement by incorporating the latest advances in multi-time-scale stochastic approximation (Han et al., 2024), leveraging the smoothness condition of the reward and transition.
We support the merit of proposed method through numerical simulations. We apply AC-SMFG to a range of environments inspired by real-world economic scenarios involving a prominent leader and a population of followers. Previous agent-based modeling (ABM) approaches to these problems simulate a large number of agents with complex interactions, leading to computational inefficiency and lack of theoretical guarantees (Evans et al., 2025). In contrast, we show that the SMFG formulation offers a more tractable alternative by summarizing the collective follower behavior through the mean field, thereby enabling faster computation of a (possibly local) optimum for the leader’s policy. Compared to the existing SMFG algorithms, AC-SMFG exhibits a significantly faster convergence rate.
@@ 1.2 Related Literature on Stackelberg Mean Field Games
(i) Continuous-time setting: (Djete, 2023) is among the first works to study SMFGs and to provide a characterization of the problem structure. Building on this, Dayanıklı and Laurière (2024) introduces a penalty-based approach that reformulates a SMFG as a single-level mean field control problem, offering a conceptual simplification. However, this approach does not come with a provably convergent learning algorithm, leaving open the question of how to efficiently and reliably solve SMFGs. (ii) Discrete-time setting: (Guo et al., 2022) consider a finite-horizon setup, and propose a minimax optimization framework (known model) for finding the policy of the leader considering a worst-case objective for the followers. Cui et al. (2024) consider finding an equilibrium in a (related, but distinct setup) of major-minor MFGs and proposes a nested-loop algorithm based on fictitious play, which approximately alternates best-response updates between the leader and the follower. Such a nested-loop structure often poses practical challenges, as it requires the inner loop (best response computation) to converge before each outer-loop update. This results in increased computational burden and reduced flexibility, particularly in large-scale or online settings where full convergence at each iteration may be infeasible or undesirable. Moreover, Cui et al. (2024) establishes only asymptotic convergence, offering limited insight into the algorithm’s performance in practice. Compared to the related literature discussed above, our paper is the first work to propose a single-loop algorithm for learning equilibria in SMFGs, supported by a non-asymptotic analysis.
@@ 3 Algorithm
The tabular softmax parameterization is considered for the purpose of mathematical analysis – optimizing over the space of softmax parameters allows us to exploit the PL structure that each player’s objective observes with respect to its parameter (Mei et al., 2020). In practice, the proposed algorithm can be applied with function approximations such as neural networks, as discussed in detail in Remark 3. We can express the policy gradients of $\omega$ and $\theta$ in the closed form below (Sutton et al., 1999).
@@
Our method exemplifies independent learning in the sense that $\omega_{k},\theta_{k},\hat{\mu}_{k}$ are each updated to optimize its own objective without any knowledge of the other two, though the samples used for the updates are from an environment determined by all variables. We note that the algorithm is single-loop and uses two trajectories of continuously generated samples [^1] The second trajectory of samples is used solely to estimate the mean field and may be eliminated in applications where there exists a generative model/oracle for the mean field., making it significantly more practical than the existing methods with complex nested loops (Dayanıklı and Laurière, 2024; Cui et al., 2024).
@@ 4 Convergence Analysis
The Lipschitz continuity is a common assumption made in the literature on MFGs (Cui and Koeppl, 2021b; Zeng et al., 2025; Cui et al., 2024). We additionally assume that they are smooth.
@@
The first part of Assumption 3 implies that the discounted occupancy measure $d_{\rho}^{\pi,\phi,\mu}$ is bounded away from zero for any $\pi,\phi,\mu$ and is a standard assumption in RL (Agarwal et al., 2021; Mei et al., 2020), while the second part can be readily satisfied by capping the norm of $\theta_{k}$ away from infinity. The assumption guarantee sufficient exploration of every state and action pair.
@@
Assumption 4 states that in a standard MFG determined by any fixed $\phi$, an “optimality-consistency operator contraction condition” holds, which implies the uniqueness of the MFE. This is a key assumption made in the existing literature on MFGs (Xie et al., 2021; Zaman et al., 2023; Yardim et al., 2023) and can be guaranteed when the regularization weight $\tau$ is large enough.
@@
Assumption 5 is again standard in analyzing stochastic approximation and RL algorithms under Markovian samples (Zou et al., 2019; Wu et al., 2020; Wang et al., 2024).
@@
Note that $\nabla_{\omega}J_{l}(\phi_{\omega},\mu)\mid_{\mu=\mu^{\star}(\phi_{\omega})}$ is a partial component of the full gradient $\nabla_{\omega}\Phi(\phi_{\omega})$. The condition states that the two gradients always make an acute angle. Without loss of generality, we assume $\eta_{1},\eta_{2}\geq 1$, as if the conditions hold with $\eta_{1},\eta_{2}<1$ they also hold with $\eta_{1},\eta_{2}=1$. Conceptually, the condition allows the leader to improve its objective by acting to the best response from the followers. Prior works on SMFGs and the closely related major-minor MFGs need to make assumptions of a similar nature, but more restrictive. For example, Cui et al. (2024) by their Assumptions 4.c) and 4.d) assumes that the leader’s reward is independent of the mean field and that the follower’s reward and transition are independent of the leader’s action. Under such conditions, the leader and follower are effectively decoupled, reducing the hierarchical structure of the problem to a near-independent setting. In contrast, with the gradient alignment assumption we permit our leader to depend on $\mu$, and follower rewards and transitions to depend on $b$. In Appendix E, we present a non-trivial MFG which satisfies the gradient alignment assumption but not the leader-follower independence condition in Cui et al. (2024).
@@ 4.1 Algorithm Complexity
As the residuals are all non-negative, Theorem 1 implies that the best iterate of Algorithm 1 converges to a stationary point of the leader’s objective with rate $\widetilde{{\cal O}}(k^{-1/2})$, where $\widetilde{{\cal O}}$ hides constants and logarithmic factors of $k$. As exactly two samples are drawn in each iteration, this translates to a sample complexity of the same order. To our knowledge, this is the first result providing a finite-time and sample complexity for learning in Stackelberg/major-minor mean field games. In contrast to the most relevant prior work (Cui et al., 2024), which establishes the asymptotic convergence of a nested-loop fictitious play method, our paper analyzes a single-loop algorithm and provides non-asymptotic guarantees. This rate improves upon that of a two-time-scale algorithm for bi-level optimization with a non-convex upper-level objective and lower-level strong convexity (Hong et al., 2023), which achieves $\widetilde{{\cal O}}(1/k^{2/5})$. Our improvement is obtained in a more challenging setting – the lower-level problem is not a strongly convex function but a MFG involving two coupled variables: the follower’s policy and the mean field. The MFG does not exhibit any convex structure, and only a non-uniform PL condition holds with respect to the follower’s policy under regularization. A key technical innovation made in our work is to handle such non-convexity, which we sketch in Remark 2, and which we believe is of independent interest and applicable to the analysis of general bi-level optimization methods with lower-level PL condition.
@@
Two techniques allow us to achieve the convergence rate of $\widetilde{{\cal O}}(k^{-1/2})$, surpassing that in Hong et al. (2023) under lower-level strong convexity.
@@
The technique highlighted above allows us to achieve the same convergence rate in SMFGs as in bi-level optimization with lower-level strong convexity. We further improve upon the rate in Hong et al. (2023), leveraging a recent advance in the multi-time-scale stochastic approximation literature. Specifically, Han et al. (2024) shows that a faster convergence rate can be achieved under lower-level strong convexity, if the lower-level learning target (equivalent to operators $\mu^{\star},\pi^{\star}$ in our context) has Lipschitz gradients. Our work adapts and extends the argument to the case where the lower-level objective is nonconvex.
@@ 5 Experiments
We conduct a comprehensive evaluation of the proposed methodology across a diverse set of canonical MFGs. Specifically, we extend three environments from MFGLib (Guo et al., 2023a), each exhibiting varying degrees of complexity. These environments, ordered by increasing difficulty, include Left-Right, Beach Bar, and Equilibrium pricing (Guo et al., 2023b), with descriptions in Appendix F. We also include in the appendix additional plots visualizing the evolution of the mean field as learning proceeds. All source code is available in the supplementary material.
Comparisons. The proposed approach is compared to the existing state-of-the-art methods for the discrete time setting: 1) a nested-loop MFG algorithm which alternates between training the leader and representative follower to convergence (OneByOne); 2) a weighted OneByOne update as presented in Algorithm 1 of (Cui et al., 2024) (CuiKoeppl); 3) a MARL implementation, where the leader and followers are updated simultaneously with PPO, as presented in (Evans et al., 2025) (ADAGE). In the MARL setting, there are two policies, one for the leader, and a shared policy (Vadori et al., 2020) for the followers, to keep the representation as fair as possible when compared to mean fields, and to prevent the need to learn $N$ independent follower policies.
@@
The first environment is a market entrance scenario inspired by the Left-Right game (Cui and Koeppl, 2021a), akin to minority games in physics and El Farol problems in complex systems, as detailed in Appendix F.1. Figure 3(a) illustrates the impact of various leader configurations, achieved by altering the leader’s objective, demonstrating how the leader learns to execute optimal actions while followers adapt effectively, steering the mean field towards the target state (see also Figure 6). The top row of Figure 3(a) highlights the leader’s influence on the resulting mean fields, while the bottom row depicts the market skew. The leader shifts the market dynamics to align with the desired behavior. In each scenario, we observe convergence towards a suitable MFE, with the MF adjusting appropriately to the task at hand. These findings underscore how follower behavior successfully adapts in response to the leader’s actions, consistently achieving appropriate equilibria.
@@
Finally, we examine an equilibrium pricing environment (Appendix F.3), as outlined in (Guo et al., 2023a, b; Cousin et al., 2011). In this environment, followers are homogeneous firms producing an identical product, with prices determined by the endogenous supply-demand equilibrium. These firms must effectively manage their inventory, production, and replenishment processes, and the leaders role is to incentivize firms to maintain a specific target inventory level. We explore various target scenarios, ranging from promoting lean operations with minimal inventories (suitable for predictable periods with stable demand) to encouraging robust firms with substantial inventory reserves (to withstand crises or demand surges). As illustrated in Figure 3(c), the leader can effectively shift the mean field of the followers to achieve the desired inventory levels. This demonstrates the leader’s ability to implement appropriate penalties or bonuses to guide followers towards either building a buffer (Figure 3(c) right) or maintaining leaner operations (Figure 3(c) left).

### v2

@@ 1 Introduction
Mean field games (MFGs) provide a framework for studying the strategic interaction among an infinite number of rational agents, with a wide range of applications in resource allocation (li2020resource), telecommunication (narasimha2019mean), and power system optimization (alasseur2020extended). An important extension of MFGs is Stackelberg mean field games (SMFGs), incorporating a hierarchical structure where a single leader agent influences a population of follower agents and enjoys a first-mover advantage. A prominent example of SMFG is optimal liquidation (almgren1997optimal), in which a large institutional investor (leader) strategically sells assets, e.g., central bank treasuries, in a market while accounting for the reactions of smaller traders (followers) who adjust their behavior, e.g., bond purchasing, in response (chen2024periodic). Other examples arise in public policy domains such as taxation, mortgage regulation, and epidemic control (zheng2020ai; mi2024learning; aurell2022optimal), where the leader, typically a government, designs interventions to promote social welfare while anticipating and influencing the collective response of the population.
Despite the increasing number of empirical studies applying the SMFG framework to real-world problems, there remains a lack of practical and theoretically grounded algorithms for learning in such hierarchical environments. The central challenge lies in the inherently coupled dynamics among the leader, a representative follower, and the mean field representing the collective behavior of the follower population. This interdependence gives rise to a complex bi-level structure: the leader’s objective depends on the equilibrium response of the follower population, which in turn is shaped by the leader’s policy. Existing approaches (dayanikli2024machine; cui2024learning) to SMFGs (and the closely related major-minor MFGs) need to make strong assumptions to decouple the leader and followers, lack non-asymptotic guarantees for finding a solution, and empirically exhibit slow convergence or tend towards sub-optimal limit points, possibly due to cyclic behaviors. In contrast, our work relaxes the assumptions and establishes a fast finite-sample rate of convergence for a simple actor-critic algorithm to the stationary point of the SMFG objective.
Figure 1: Theoretical convergence rates. The proposed algorithm is the first to have non-asymptotic convergence guarantees (black). Existing algorithms in similar settings are only known to converge asymptotically (red). Convergence rate of a single-loop algorithm for bi-level optimization (hong2023two) under a non-convex upper-level objective and lower-level strong convexity (blue).
@@ 1.1 Main Contributions
Notably, our convergence rate surpasses that of a comparable single-loop algorithm for bi-level optimization under lower-level strong convexity (hong2023two) ($\widetilde{{\cal O}}(k^{-1/2})$ vs $\widetilde{{\cal O}}(k^{-2/5})$). The policy optimization problem in a SMFG can be viewed as bi-level optimization, where the upper-level objective is non-convex and the lower-level problem is to solve a standard MFG that does not exhibit any convexity. The reason that we achieve a better rate in a more challenging setting is two-fold. First, as a key technical innovation, we develop an analytical argument to handle bi-level optimization with lower-level Polyak-Lojasiewicz (PL) condition, which the MFG objective satisfies with respect to the follower’s softmax policy parameter under regularization (mei2020global). This technique allows us to obtain the same convergence rate in SMFGs as in bi-level optimization with lower-level strong convexity, and may be of independent interest and applicable to general bi-level optimization methods. Second, we make the rate improvement by incorporating the latest advances in multi-time-scale stochastic approximation (han2024finite), leveraging the smoothness condition of the reward and transition.
We support the merit of proposed method through numerical simulations. We apply AC-SMFG to a range of environments inspired by real-world economic scenarios involving a prominent leader and a population of followers. Previous agent-based modeling (ABM) approaches to these problems simulate a large number of agents with complex interactions, leading to computational inefficiency and lack of theoretical guarantees (evans2025adage). In contrast, we show that the SMFG formulation offers a more tractable alternative by summarizing the collective follower behavior through the mean field, thereby enabling faster computation of a (possibly local) optimum for the leader’s policy. Compared to the existing SMFG algorithms, AC-SMFG exhibits a significantly faster convergence rate.
@@ 1.2 Related Literature on Stackelberg Mean Field Games
(i) Continuous-time setting: (djete2023stackelberg) is among the first works to study SMFGs and to provide a characterization of the problem structure. Building on this, dayanikli2024machine introduces a penalty-based approach that reformulates a SMFG as a single-level mean field control problem, offering a conceptual simplification. However, this approach does not come with a provably convergent learning algorithm, leaving open the question of how to efficiently and reliably solve SMFGs. (ii) Discrete-time setting: (guo2022optimization) consider a finite-horizon setup, and propose a minimax optimization framework (known model) for finding the policy of the leader considering a worst-case objective for the followers. cui2024learning consider finding an equilibrium in a (related, but distinct setup) of major-minor MFGs and proposes a nested-loop algorithm based on fictitious play, which approximately alternates best-response updates between the leader and the follower. Such a nested-loop structure often poses practical challenges, as it requires the inner loop (best response computation) to converge before each outer-loop update. This results in increased computational burden and reduced flexibility, particularly in large-scale or online settings where full convergence at each iteration may be infeasible or undesirable. Moreover, cui2024learning establishes only asymptotic convergence, offering limited insight into the algorithm’s performance in practice. Compared to the related literature discussed above, our paper is the first work to propose a single-loop algorithm for learning equilibria in SMFGs, supported by a non-asymptotic analysis.
@@ 3 Algorithm
The tabular softmax parameterization is considered for the purpose of mathematical analysis – optimizing over the space of softmax parameters allows us to exploit the PL structure that each player’s objective observes with respect to its parameter (mei2020global). In practice, the proposed algorithm can be applied with function approximations such as neural networks, as discussed in detail in Remark 3. We can express the policy gradients of $\omega$ and $\theta$ in the closed form below (sutton1999policy).
@@
Our method exemplifies independent learning in the sense that $\omega_{k},\theta_{k},\hat{\mu}_{k}$ are each updated to optimize its own objective without any knowledge of the other two, though the samples used for the updates are from an environment determined by all variables. We note that the algorithm is single-loop and uses two trajectories of continuously generated samples [^1] The second trajectory of samples is used solely to estimate the mean field and may be eliminated in applications where there exists a generative model/oracle for the mean field., making it significantly more practical than the existing methods with complex nested loops (dayanikli2024machine; cui2024learning).
@@ 4 Convergence Analysis
The Lipschitz continuity is a common assumption made in the literature on MFGs (cui2021learning; zeng2024learning; cui2024learning). We additionally assume that they are smooth.
@@
The first part of Assumption 3 implies that the discounted occupancy measure $d_{\rho}^{\pi,\phi,\mu}$ is bounded away from zero for any $\pi,\phi,\mu$ and is a standard assumption in RL (agarwal2021theory; mei2020global), while the second part can be readily satisfied by capping the norm of $\theta_{k}$ away from infinity. The assumption guarantee sufficient exploration of every state and action pair.
@@
Assumption 4 states that in a standard MFG determined by any fixed $\phi$, an “optimality-consistency operator contraction condition” holds, which implies the uniqueness of the MFE. This is a key assumption made in the existing literature on MFGs (xie2021learning; zaman2023oracle; yardim2023policy) and can be guaranteed when the regularization weight $\tau$ is large enough.
@@
Assumption 5 is again standard in analyzing stochastic approximation and RL algorithms under Markovian samples (zou2019finite; wu2020finite; wang2024non).
@@
Note that $\nabla_{\omega}J_{l}(\phi_{\omega},\mu)\mid_{\mu=\mu^{\star}(\phi_{\omega})}$ is a partial component of the full gradient $\nabla_{\omega}\Phi(\phi_{\omega})$. The condition states that the two gradients always make an acute angle. Without loss of generality, we assume $\eta_{1},\eta_{2}\geq 1$, as if the conditions hold with $\eta_{1},\eta_{2}<1$ they also hold with $\eta_{1},\eta_{2}=1$. Conceptually, the condition allows the leader to improve its objective by acting to the best response from the followers. Prior works on SMFGs and the closely related major-minor MFGs need to make assumptions of a similar nature, but more restrictive. For example, cui2024learning by their Assumptions 4.c) and 4.d) assumes that the leader’s reward is independent of the mean field and that the follower’s reward and transition are independent of the leader’s action. Under such conditions, the leader and follower are effectively decoupled, reducing the hierarchical structure of the problem to a near-independent setting. In contrast, with the gradient alignment assumption we permit our leader to depend on $\mu$, and follower rewards and transitions to depend on $b$. In Appendix E, we present a non-trivial MFG which satisfies the gradient alignment assumption but not the leader-follower independence condition in cui2024learning.
@@ 4.1 Algorithm Complexity
As the residuals are all non-negative, Theorem 1 implies that the best iterate of Algorithm 1 converges to a stationary point of the leader’s objective with rate $\widetilde{{\cal O}}(k^{-1/2})$, where $\widetilde{{\cal O}}$ hides constants and logarithmic factors of $k$. As exactly two samples are drawn in each iteration, this translates to a sample complexity of the same order. To our knowledge, this is the first result providing a finite-time and sample complexity for learning in Stackelberg/major-minor mean field games. In contrast to the most relevant prior work (cui2024learning), which establishes the asymptotic convergence of a nested-loop fictitious play method, our paper analyzes a single-loop algorithm and provides non-asymptotic guarantees. This rate improves upon that of a two-time-scale algorithm for bi-level optimization with a non-convex upper-level objective and lower-level strong convexity (hong2023two), which achieves $\widetilde{{\cal O}}(1/k^{2/5})$. Our improvement is obtained in a more challenging setting – the lower-level problem is not a strongly convex function but a MFG involving two coupled variables: the follower’s policy and the mean field. The MFG does not exhibit any convex structure, and only a non-uniform PL condition holds with respect to the follower’s policy under regularization. A key technical innovation made in our work is to handle such non-convexity, which we sketch in Remark 2, and which we believe is of independent interest and applicable to the analysis of general bi-level optimization methods with lower-level PL condition.
@@
Two techniques allow us to achieve the convergence rate of $\widetilde{{\cal O}}(k^{-1/2})$, surpassing that in hong2023two under lower-level strong convexity.
@@
The technique highlighted above allows us to achieve the same convergence rate in SMFGs as in bi-level optimization with lower-level strong convexity. We further improve upon the rate in hong2023two, leveraging a recent advance in the multi-time-scale stochastic approximation literature. Specifically, han2024finite shows that a faster convergence rate can be achieved under lower-level strong convexity, if the lower-level learning target (equivalent to operators $\mu^{\star},\pi^{\star}$ in our context) has Lipschitz gradients. Our work adapts and extends the argument to the case where the lower-level objective is nonconvex.
@@ 5 Experiments
We conduct a comprehensive evaluation of the proposed methodology across a diverse set of canonical MFGs. Specifically, we extend three environments from MFGLib (mfglib), each exhibiting varying degrees of complexity. These environments, ordered by increasing difficulty, include Left-Right, Beach Bar, and Equilibrium pricing (guo2023general), with descriptions in Appendix F. We also include in the appendix additional plots visualizing the evolution of the mean field as learning proceeds. All source code is available in the supplementary material.
Comparisons. The proposed approach is compared to the existing state-of-the-art methods for the discrete time setting: 1) a nested-loop MFG algorithm which alternates between training the leader and representative follower to convergence (OneByOne); 2) a weighted OneByOne update as presented in Algorithm 1 of (cui2024learning) (CuiKoeppl); 3) a MARL implementation, where the leader and followers are updated simultaneously with PPO, as presented in (evans2025adage) (ADAGE). In the MARL setting, there are two policies, one for the leader, and a shared policy (vadori2020calibration) for the followers, to keep the representation as fair as possible when compared to mean fields, and to prevent the need to learn $N$ independent follower policies.
@@
The first environment is a market entrance scenario inspired by the Left-Right game (cui2021approximately), akin to minority games in physics and El Farol problems in complex systems, as detailed in Appendix F.1. Figure 3(a) illustrates the impact of various leader configurations, achieved by altering the leader’s objective, demonstrating how the leader learns to execute optimal actions while followers adapt effectively, steering the mean field towards the target state (see also Figure 6). The top row of Figure 3(a) highlights the leader’s influence on the resulting mean fields, while the bottom row depicts the market skew. The leader shifts the market dynamics to align with the desired behavior. In each scenario, we observe convergence towards a suitable MFE, with the MF adjusting appropriately to the task at hand. These findings underscore how follower behavior successfully adapts in response to the leader’s actions, consistently achieving appropriate equilibria.
@@
Finally, we examine an equilibrium pricing environment (Appendix F.3), as outlined in (mfglib; guo2023general; cousin2011mean). In this environment, followers are homogeneous firms producing an identical product, with prices determined by the endogenous supply-demand equilibrium. These firms must effectively manage their inventory, production, and replenishment processes, and the leaders role is to incentivize firms to maintain a specific target inventory level. We explore various target scenarios, ranging from promoting lean operations with minimal inventories (suitable for predictable periods with stable demand) to encouraging robust firms with substantial inventory reserves (to withstand crises or demand surges). As illustrated in Figure 3(c), the leader can effectively shift the mean field of the followers to achieve the desired inventory levels. This demonstrates the leader’s ability to implement appropriate penalties or bonuses to guide followers towards either building a buffer (Figure 3(c) right) or maintaining leaner operations (Figure 3(c) left).